# A Competitive Showcase of Quantum versus Classical Algorithms in Energy Coalition Formation

### v1

@@ III Results
Randomly generated problems: To enhance the robustness of our findings and gain deeper insights into the performance scalability of D-Wave, we validate our results on larger-scale random problems by generating our problem instances randomly. Solving power flow equations for more than a few tens of prosumers demands substantial computational resources, and for the scope of this work, random generation suffices. For smaller numbers of prosumers, the distribution of generated problems, after transforming from CFG to ISG, follows a Gaussian distribution centered around zero with a width of 0.2 across all prosumer sizes. Therefore, for larger sizes, we generate our problem instances using the same Gaussian distribution but intentionally vary the width to 2. This deliberate variation allows us to investigate the correlation between solvers’ performance and distribution characteristics, which may align with different choices of constraints and parameters in the problem formulation
The outcome of this investigation is presented in Fig. 3. For a more comprehensive benchmark, we additionally employ two other classical solvers: simulated annealing and Tabu search. Given that the exact solution is unknown for large system sizes, we assess the quality of solutions by calculating the sum of values from the optimum coalitions, divided by the number of prosumers. A higher value signifies a superior solution. Dwave consistently demonstrates performance in line with the results depicted in Fig. 2 and competes effectively with the classical solvers. Up to the scale implemented on the hardware, Dwave demonstrates a more favorable runtime scaling as a function of problem size, as depicted in Fig. 3(b). The runtime of QB-solve, Tabu search, and simulated annealing scales cubically with varying prefactors, with Tabu search exhibiting the lowest prefactor. In contrast, Dwave’s runtime scales as $\sqrt{n}$ with $n$ the number of prosumers. At the same time, all solvers produce solutions of almost equal values, as shown in Fig. 3(a).
@@
When using the D-Wave quantum computer to solve a problem, the connectivity of the logical qubits in the problem must be mapped onto the hardware’s connectivity graph, which, for Advantage quantum processing units, is based on the Pegasus graph topology. This mapping process typically necessitates additional qubits, known as physical qubits, beyond those required by the original problem. Figure 3 (c) illustrates the scaling of physical qubits with the number of logical qubits (prosumers).
@@ IV Conclusion and outlook
The formation of energy communities signifies a pivotal shift towards decentralized and sustainable energy management. We employ the characteristic function game, offering a framework to comprehend how prosumers can unite in coalitions to maximize their collective advantages. A pivotal aspect of our approach is the exploration of optimal coalition compositions, balancing the benefits derived from increased resources for cooperative optimization against the network constraints and losses incurred when coalitions are dispersed. We conduct a benchmark comparing classical solvers with quantum solvers in this problem domain. Our findings showcase that the Dwave machine outperforms QAOA across hardware in terms of solution quality. However, it is important to note that the implementation of QAOA lacks intensive optimization in hardware resources and error mitigation schemes and there is still significant potential for further improvement in the performance of QAOA. Interestingly, QAOA demonstrates comparable runtime scaling with Dwave, albeit with a notably larger prefactor. Dwave exhibits competitive performance compared to our approximate classical solvers. While it achieves solutions of equal quality to those found by Tabu search and SA, its runtime is notably more favorable. This parity in performance highlights the promising potential of quantum algorithms in tackling optimization problems. The competitive performance shown by Dwave on our fully connected problems stands in contrast to its poor performance in previous studies Mohseni et al. (2022); Au-Yeung et al. (2023); Vert et al. (2021); Carugno et al. (2022).
It is essential to acknowledge that classical solvers have benefited from decades of development, while quantum solvers are still in their early stages of evolution. Therefore, the observed competitive performance highlights the evolving potential of quantum computing in optimization problems.
@@ Acknowledgements.
N.M. would like to thank Supreeth Mysore, Ivan Angelov, Stefan Wörner, and Gabriele Agliardi for fruitful discussions. T. M. acknowledges the UK Engineering and Physical Sciences Research Council (EPSRC) (grant reference number EP/Y004418/1). This work was supported by the German Federal Ministry of Education and Research under the funding program “Förderprogramm Quantentechnologien – von den Grundlagen zum Markt” (funding program quantum technologies — from basic research to market), project Q-Grid, 13N16177.
@@ I.1.1 Quantum Annealing
Quantum annealing (QA) is a heuristic algorithm based on the quantum adiabatic theorem Farhi et al. (2000); Albash and Lidar (2018b). In this algorithm, the system is initially prepared in the ground state of a Hamiltonian $H_{0}$ where its ground state is known. A common choice for this initial Hamiltonian is $H_{0}=-\sum_{i=1}^{N}\sigma_{i}^{x}$, where $N$ denotes the number of qubits, and the ground state is the uniform superposition of all possible configurations $|+\rangle^{\otimes N}$ with $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$. The Hamiltonian is gradually reweighted to the desired problem Hamiltonian $H_{C}$ according to
@@
where $\lambda(t)\in[0,1]$ is the annealing schedule. The annealing process can be viewed as $H_{0}$ introducing quantum fluctuations originating from the non-commutability of $H_{C}$ and $H_{0}$. These fluctuations are gradually reduced to reach the low-energy configuration of the classical energy function $H_{C}$. Based on the quantum adiabatic theorem, if one performs the sweep sufficiently slowly, the system remains in its instantaneous ground state throughout the evolution (Fig. S1 (a) ) Farhi et al. (2000); Albash and Lidar (2018b). $H_{C}$ in our formulation is Eq. (6) at each min cut iteration. The number of qubits is then the number of prosumers. The use of quantum fluctuations in QA has been hypothesized as a potential resource for a speedup over classical methods. Quantum tunneling allows the system to pass through energy barriers that have a higher energy than available in the state Denchev et al. (2016).
@@ I.2.3 Simulated annealing
I.2.4 Random Sampler
@@ II Technical details on problem formulation
    $-p^{\text{initial}}_{it}(1+\epsilon)\leq p_{i,t}\leq p^{\text{initial}}_{it}(1+\epsilon)$

### v2

@@ III Results
Randomly generated problems: To enhance the robustness of our findings and gain deeper insights into the performance scalability of DWave, we validate our results on larger-scale random problems by generating our problem instances randomly. Solving power flow equations for more than a few tens of prosumers (required to generate problem instances) demands substantial computational resources, and for the scope of this work, random generation suffices. For smaller numbers of prosumers, the distribution of generated problems, after transforming from CSG to ISG, follows a Gaussian distribution centered around zero with a width of 0.2 across all prosumer sizes. Therefore, for larger sizes (above 22 prosumers), we generate our problem instances using the same Gaussian distribution but intentionally vary the width to 2. This deliberate variation allows us to investigate the correlation between solvers’ performance and distribution characteristics, which may align with different choices of constraints and parameters in the problem formulation
The outcome of this investigation is presented in Fig. 3. For a more comprehensive benchmark, we additionally employ two other classical solvers: simulated annealing and Tabu search. For instances with more than 50 prosumers, Gurobi fails to find a provably optimal solution and fails to terminate even after several hours of computation. Additionally, we observed that the objective value no longer improves significantly after some time. Therefore, we introduced a termination criterion: the optimization process is stopped when the change in the objective value is less than 1e-8, and no significant improvement is observed within a predefined time window of 1 second. Note that if no improvement occurs within this 1-second window, the time is not included in the reported runtime. We consider the best solution found by Gurobi as the reference, even though its optimality remains unproven and other solvers may find a better solution.
The optimality gap—the difference between the best solution found by the solver and the upper bound of the optimal solution, derived from the dual problem—varies across different problem instances and iterations, particularly during the process of splitting coalitions into bipartite groups. We report the optimality gap for the first iteration, namely the initial split of the grand coalition into two partitions. Averaged across all problem instances, the gap is $17\%,26\%,25\%,29\%,28\%,30\%,29\%$ for prosumer sizes of 50, 100, 150, 200, 250, 300 and 350, respectively. In most cases, the optimality gap reaches zero in later iterations.
DWave consistently demonstrates performance in line with the results depicted in Fig. 2 and competes effectively with the classical solvers. In terms of solution quality, for problem sizes up to 150 prosumers, the largest scale tested on the hardware, D-Wave consistently finds solutions similar to those obtained by Gurobi, except for the case with 150 prosumers, where D-Wave’s solution is $1\%$ suboptimal. Within this range of prosumers, Tabu search consistently finds solutions that are either slightly better or worse than those found by Gurobi, within less than $1\%$ difference. For instances with over 200 prosumers, where no results are available from D-Wave, Gurobi consistently identifies the best solution among classical solvers, and all solvers outperform SA.
In terms of run time, DWave demonstrates a more favorable runtime scaling as a function of problem size, as depicted in Fig. 3(b). The runtime of Gurobi, QB-solve, Tabu search, and simulated annealing scales cubically with varying prefactors, with Tabu search exhibiting the lowest prefactor. In contrast, DWave’s runtime scales as $\sqrt{n}$ with $n$ the number of prosumers.
@@
When using the DWave quantum computer to solve a problem, the connectivity of the logical qubits in the problem must be mapped onto the hardware’s connectivity graph, which, for Advantage quantum processing units, is based on the Pegasus graph topology. This mapping process typically necessitates additional qubits, known as physical qubits, beyond those required by the original problem. Figure 3 (c) illustrates the scaling of physical qubits with the number of logical qubits (prosumers).
In Fig. 3 (d), we present the energy distribution (for a randomly generated instance with 100 prosumers) sampled from QAOA enhanced with Fire Opal’s error suppression pipeline Sachdeva et al. (2024), DWave, and a random sampler. These distributions correspond to the first split of the grand coalition into two partitions. The vertical cyan line marks the solution found by Gurobi, though its optimality is unproven. Notably, DWave recovers the same solution, while the QAOA distribution closely resembles that of the random sampler. The circuit depth for QAOA in this experiment applying the Fireopal’s transpiler is 11k. Note that no postprocessing is applied to the QAOA output.
@@ IV Conclusion and outlook
The formation of energy communities signifies a pivotal shift towards decentralized and sustainable energy management. In this work, we propose a novel formulation of energy coalition formation that can be encapsulated within the framework of Coalition Structure Generation to comprehend how prosumers can unite in coalitions to maximize their collective advantages. We also introduce a novel method for transforming this formulation into a QUBO formulation for quantum solvers. We then conduct a benchmark comparing classical solvers with quantum solvers in this problem domain. Our findings showcase that the DWave machine outperforms 1-round QAOA on IBM hardware in terms of solution quality. D-Wave exhibits more favorable time scaling with problem size than classical approximate solvers. However, for the system sizes we evaluated, it remains slower than some classical methods, such as QB-Solve. In terms of solution quality, D-Wave matches our best classical solvers (e.g., Gurobi and Tabu Search) for up to 100 prosumers (qubits) and achieves solutions within a $1\%$ optimality gap from Gurobi for problems with 150 prosumers (150 logical qubits).
@@ Acknowledgements.
N.M. and C.O. would like to thank Supreeth Mysore, Ivan Angelov, Stefan Wörner, Antonio Macaluso, Gabriele Agliardi, and Rowen Wu for fruitful discussions. T. M. acknowledges the UK Engineering and Physical Sciences Research Council (EPSRC) (grant reference number EP/Y004418/1). This work was supported by the German Federal Ministry of Education and Research under the funding program “Förderprogramm Quantentechnologien – von den Grundlagen zum Markt” (funding program quantum technologies — from basic research to market), project Q-Grid, 13N16177.
@@ I.1.1 Quantum Annealing
Quantum annealing (QA) is a heuristic algorithm based on the quantum adiabatic theorem Farhi et al. (2000); Albash and Lidar (2018). In this algorithm, the system is initially prepared in the ground state of a Hamiltonian $H_{0}$ where its ground state is known. A common choice for this initial Hamiltonian is $H_{0}=-\sum_{i=1}^{N}\sigma_{i}^{x}$, where $N$ denotes the number of qubits, and the ground state is the uniform superposition of all possible configurations $|+\rangle^{\otimes N}$ with $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$. The Hamiltonian is gradually reweighted to the desired problem Hamiltonian $H_{C}$ according to
@@
where $\lambda(t)\in[0,1]$ is the annealing schedule. The annealing process can be viewed as $H_{0}$ introducing quantum fluctuations originating from the non-commutability of $H_{C}$ and $H_{0}$. These fluctuations are gradually reduced to reach the low-energy configuration of the classical energy function $H_{C}$. Based on the quantum adiabatic theorem, if one performs the sweep sufficiently slowly, the system remains in its instantaneous ground state throughout the evolution (Fig. S1 (a) ) Farhi et al. (2000); Albash and Lidar (2018). $H_{C}$ in our formulation is Eq. (6) at each min cut iteration. The number of qubits is then the number of prosumers. The use of quantum fluctuations in QA has been hypothesized as a potential resource for a speedup over classical methods. Quantum tunneling allows the system to pass through energy barriers that have a higher energy than available in the state Denchev et al. (2016).
@@ I.2.3 Simulated annealing
I.2.4 Gurobi
Gurobi Gurobi Optimization, LLC (2024) Gurobi is a state-of-the-art commercial solver for a wide range of optimization problems, including linear programming, mixed-integer programming, and quadratic programming. It uses advanced techniques like branch-and-cut, combining the cutting plane method to tighten the solution space and the branch-and-bound algorithm to systematically explore possible solutions. In Gurobi, the optimality gap is defined as the difference between the best solution found by the solver, and the best upper bound estimation of the optimal solution, estimated through the dual problem. A zero gap implies that Gurobi’s solution is exact.
I.2.5 Random Sampler
@@ II Technical details on problem formulation
    $p^{\text{initial}}_{it}(1+\epsilon)\leq p_{i,t}\leq p^{\text{initial}}_{it}(1+\epsilon)$